I have to write a two-step equation.
A rectangle has a width of 6 inches and a perimeter of 26 inches. What is the length of the rectangle

Answer by guitarheroguy(26)

OK, well first we have to figure out both sides of the width
width=6 inches
6*2=12
Total perimiter=26
perimeter-width=length
26-12=14
14/2=7
The width is 7
Let's Check!
7+7+6+6=26